OpenStudy (anonymous):

Determine wheter the function is one-on-one. If it is, find a formula for its inverse. f(x)=4/5x+a

OpenStudy (anonymous):

is this \[f(x)=\frac{4}{5x+a}\]?

OpenStudy (anonymous):

if so, it is indeed one to one. you can find the inverse by writing \(x=\frac{4}{5y+a}\) and solving for \(y\)

OpenStudy (anonymous):

\[x=\frac{4}{5y+a}\] \[5xy+ax=4\] \[5xy=4-ax\] \[y=\frac{4-ax}{5x}\] and so \[f^{-1}(x)=\frac{4-ax}{5x}\]